OpenStudy (anonymous):

I'll put the bases into () . Solve log(1-x)3-log(1-x)2-0,5=0

OpenStudy (anonymous):

log(1-x)3 - log(1-x)2 - 0.5 = 0 log(1-x)3/2 = 0.5 : : Combine log functions using log properties and move 0.5 to other side (1-x)^0.5 = 3/2 : : (1-x) is the base with the exponent being 0.5 (1-x) = (3/2)^2 : : 0.5 exponent is the same as square rooting it x = 1 - (3/2)^2

OpenStudy (anonymous):

Thank you !

OpenStudy (anonymous):

no problem

OpenStudy (anonymous):

I have on more problem , if you have time .

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

Log(x)log(2x)2=log(4x)2

OpenStudy (anonymous):

And also , do you have any idea on how to solve (1-x)^4

OpenStudy (anonymous):

if i knew what is on the other side of the equation... that other log problem is tricky

OpenStudy (anonymous):

Other side is x^2

OpenStudy (anonymous):

(1-x)^4=x^2

OpenStudy (anonymous):

((1-x)^2)^2 = x^2 (1-x)^2 = x 1-x = +- x ^(1/2) 0 = +- x^(1/2) + x - 1 ... i got stuck here

OpenStudy (anonymous):

Okay , but what about that first one?

OpenStudy (anonymous):

that one involves a property of logs that i can't remember... sorry about that. Try going over the properties of logs

OpenStudy (anonymous):

Okay , thank you so much for your time !